A manufacturer of electronic components is interested in determining the lifetime of a certain type of battery. A sample, in hours of life, is as follows




123 116 122 110 175 126 125 111 118 117




(i) Determine the sample mean and median




(ii) What feature in this data set is responsible for the substantial difference between the two?





1
Expert's answer
2022-04-27T12:24:55-0400

(i)

Sorted Data Set: 110, 111, 116, 117, 118, 122, 123, 125, 126, 175


"mean=\dfrac{1}{10}(110+111+116+117"

"+118+122+123+125+126+175)=124.3"


"median=\dfrac{118+122}{2}=120"

(ii)

The extreme value (175 hours) is responsible  for the substantial difference between the sample mean and sample median.